Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability Approaches

Expert replies
by tnkippen » Tue Nov 13, 2012 7:38 am
I have been trying to use combinations more to solve probability questions, but I am finding a discrepancy between the combinations approach and the traditional approach on the problem below.

"There are 6 balls: A, B, C, D, E, F. What is the probability that a package containing two different types of balls contains E?"

Using combinations, I calculated: (5C5)(5C1)/(6C2) = 5/15 = 1/3

But using P(A or B) = P(A) + P(B) - P(A and B), I got:
P(A) = 1/6, P(B) = 1/5, P(A and B) = 0, so:
(1/6) + (1/5) = 11/30

What am I doing wrong above?
Join the discussion
Source: — Problem Solving |

by Brent@GMATPrepNow » Tue Nov 13, 2012 8:08 am
tnkippen wrote:I have been trying to use combinations more to solve probability questions, but I am finding a discrepancy between the combinations approach and the traditional approach on the problem below.

"There are 6 balls: A, B, C, D, E, F. What is the probability that a package containing two different types of balls contains E?"

Using combinations, I calculated: (5C5)(5C1)/(6C2) = 5/15 = 1/3

But using P(A or B) = P(A) + P(B) - P(A and B), I got:
P(A) = 1/6, P(B) = 1/5, P(A and B) = 0, so:
(1/6) + (1/5) = 11/30

What am I doing wrong above?
The problem is highlighted above (in green)

P(E is 1 of the 2 balls) = P(E selected first OR E selected second)
= P(E selected first) + P(E selected second) - P(E selected first and second)
= P(E selected first) + P(E selected second) - 0 (since E cannot be selected 1st and 2nd)

P(E selected first) = P(E selected 1st AND some other letter selected 2nd)
= P(E selected 1st) x P(some other letter selected 2nd)
= (1/6)(5/5)
= 1/6

P(E selected second) = P(some other letter selected 1st AND E selected 2nd)
= P(some other letter selected 1st) x P(E selected 2nd)
= (5/6)(1/5)
= 1/6

So, P(E is 1 of the 2 balls) = 1/6 + 1/6 = 1/3

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by tnkippen » Tue Nov 13, 2012 8:15 pm
Thanks again, Brent. With the combinations approach, why don't I need to multiply (5C5)(5C1)/(6C2) by 2 (to represent the chance that the E could occur first or second)?
Join the discussion

by sana.noor » Sun Oct 27, 2013 11:39 pm
we can select 2 balls from 6 balls as 6C2 = 15 ways
now withdraw E, and rest we are left with 5 balls. drawing one ball from rest of the five balls 5C1 = 5 ways. now E is already selected. 5/15 = 1/3 ways we can combine one ball out of 5 different balls with E.
Isnt it too easy?
Work hard in Silence, Let Success make the noise.

If you found my Post really helpful, then don't forget to click the Thank/follow me button. :)
Join the discussion

by GMATGuruNY » Mon Oct 28, 2013 2:52 am
There are 6 balls: A, B, C, D, E, F. What is the probability that a package containing two different types of balls contains E?
Alternate approach:

P(E is selected) = 1 - P(E is NOT selected).

P(E is not selected):
P(1st ball is not E) = 5/6. (Of the 6 balls, 5 are not E.)
P(2nd ball is not E) = 4/5. (Of the 5 remaining balls, 4 are not E.)
Since both events must happen for E not to be selected, we MULTIPLY the probabilities:
5/6 * 4/5 = 2/3.

Thus:
P(E is selected) = 1 - 2/3 = 1/3.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion